# When Fusion Helps and When It Breaks: View-Aligned Robustness in Same-Source Financial Imaging

### v1

@@ Appendix A Implementation Details for Input Parameterization
Table 5 summarizes the input parameterization and normalization settings.
Raw input.
@@
Optional per-channel standardization.
@@
Valid-range clipping.
@@
Threat model defined in raw space.
The threat model is always defined on $\mathbf{x}_{\mathrm{raw}}$ with an $\ell_{\infty}$ constraint $\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon$, where $\|\boldsymbol{\delta}\|_{\infty}=\max_{i}|\delta_{i}|$. When the model consumes standardized inputs, we enforce the same underlying raw-space budget by converting
$\epsilon_{\mathrm{norm}}=\epsilon\oslash\boldsymbol{\sigma},\qquad\alpha_{\mathrm{norm}}=\alpha\oslash\boldsymbol{\sigma},$ (14)
@@
Item Setting
Image storage 8-bit grayscale per view
Load / scaling ToTensor() $\Rightarrow\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$
Normalization (use_norm) $\mathbf{x}_{\mathrm{norm}}=(\mathbf{x}_{\mathrm{raw}}-\mu)\oslash\sigma$
$\mu,\sigma$ estimation Randomly sample $\min(n_{\text{norm}},|{\rm train}|)$ training images without replacement (default $n_{\text{norm}}{=}512$); compute per-channel mean/std over $N{\times}H{\times}W$; clamp $\sigma\leftarrow\max(\sigma,10^{-6})$
Clipping (raw) $[0,1]$
Clipping (norm) $\big[(0-\mu)\oslash\sigma,(1-\mu)\oslash\sigma\big]$
Threat model space raw space; when attacking norm, use Eq. 14
Appendix B Model Architecture and Fusion Details
Notation.
@@ Appendix B Model Architecture and Fusion Details
Lite-CNN.
Our Lite-CNN baseline is a compact CNN with three convolution blocks and two pooling operations, followed by a 2-layer MLP head. Concretely: (1) $\mathrm{Conv}_{3\times 3}(C\!\rightarrow\!16)$ + ReLU; (2) $\mathrm{Conv}_{3\times 3}(16\!\rightarrow\!16)$ + ReLU + $\mathrm{MaxPool}(2,2)$; (3) $\mathrm{Conv}_{3\times 3}(16\!\rightarrow\!32)$ + ReLU + $\mathrm{MaxPool}(3,3)$; then flatten and an MLP head $\mathrm{Linear}(\cdot\!\rightarrow\!256)\!\rightarrow\!\mathrm{Dropout}\!\rightarrow\!\mathrm{Linear}(256\!\rightarrow\!128)\!\rightarrow\!\mathrm{Linear}(128\!\rightarrow\!2)$. The flatten dimension is inferred automatically from the input resolution ($\times256$).
ResNet18-P.
@@
Late fusion with optional consistency (*-Cons).
@@
Early fusion two-view (C=2C{=}2) Same network with C=2C{=}2 at input (channel stacking) Same network with C=2C{=}2 at input (adapted first conv)
Late fusion encoders (*-Cons) Two Lite-CNN trunks (C=1C{=}1 each), output d=128d{=}128 Two ResNet-18 trunks (C=1C{=}1 each), output d=512d{=}512
Fused head (*-Cons) $(256\!\rightarrow\!2) Dropout + Linear$$(1024\!\rightarrow\!2)$
@@
Appendix C Additional Evaluation Details
From probabilities to labels.
Reporting.
@@ Appendix C Additional Evaluation Details
Appendix D Attack Hyperparameters and Implementation Notes
Perturbation budgets.
Unless stated otherwise, perturbation budgets are defined in raw $[0,1]$ pixel space under an $\ell_{\infty}$ constraint. We follow the main text and evaluate $\epsilon\in E$ defined in Eq. (4), where $\epsilon=0$ corresponds to clean evaluation. When sweeping budgets, we run the evaluation once per $\epsilon$. Table 7 summarizes the attack settings.
Loss and attack objective.
@@ Appendix D Attack Hyperparameters and Implementation Notes
$\max_{\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon}\ \mathcal{L}_{\mathrm{CE}}\!\left(f(\mathbf{x}+\boldsymbol{\delta}),y\right).$
@@
FGSM (one-step).
FGSM is implemented as a single-step update followed by clipping to the valid input range.
PGD (multi-step).
PGD performs $K=10$ iterative updates with step size $\alpha=\epsilon/K$ and projects after each step to enforce the $\ell_{\infty}$ constraint. We disable random initialization within the $\ell_{\infty}$ ball (random_start=False).
Table 7: Attack settings used in our experiments. Budgets are defined in raw $[0,1]$ space. When attacking standardized inputs, we convert $(\epsilon,\alpha)$ using Eq. 14.
Attack Objective Budget Steps $K$ Step size $\alpha$ Random init.
FGSM untargeted CE (maximize) $\epsilon\in E$ 1 $\alpha=\epsilon$ No
PGD untargeted CE (maximize) $\epsilon\in E$ 10 $\alpha=\epsilon/10$ No
Appendix E Additional Stress Test at $\epsilon_{\mathrm{adv}}=2/255$
Why include $2/255$.
We exclude $2/255$ from the main budget grid to focus on imperceptible perturbations, but report it here as an additional stress test.
Key observation.
@@ Appendix E Additional Stress Test at $\epsilon_{\mathrm{adv}}=2/255$
Table 8: PGD robustness stress test including $\epsilon_{\mathrm{adv}}=2/255$. We report MCC at two representative budgets ($0.25/255$ and $2/255$) and the worst-case drop $\Delta_{\mathrm{worst}}$ over nonzero budgets (Eq. (9)).
Model Input Attack (PGD) MCC@0 MCC@0.25/255 MCC@2/255 $\boldsymbol{\Delta_{\mathrm{worst}}}$
Lite-CNN ohlcv on ohlcv 0.109 0.054 -0.271 -0.380
Lite-CNN indic on indic 0.079 0.024 -0.450 -0.529
Lite-CNN both (early) on ohlcv (ch0) 0.089 0.030 -0.318 -0.407
Lite-CNN both (early) on indic (ch1) 0.089 0.033 -0.302 -0.391
Lite-CNN both (early) on both 0.089 -0.014 -0.611 -0.700
Lite-CNN-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.121 0.081 -0.199 -0.319
Lite-CNN-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.121 0.113 0.048 -0.073
Lite-CNN-Cons ($\lambda{=}1$) both (late) on both 0.121 0.071 -0.278 -0.398
ResNet18-P ohlcv on ohlcv 0.019 -0.443 -0.996 -1.015
ResNet18-P indic on indic 0.082 -0.999 -1.000 -1.082
ResNet18-P-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.095 -0.211 -0.961 -1.056
ResNet18-P-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.095 -0.959 -1.000 -1.095
ResNet18-P-Cons ($\lambda{=}1$) both (late) on both 0.095 -0.983 -1.000 -1.095
Appendix F Full Branch Diagnostics for Late-Fusion Models
This section reports the full branch-level diagnostics for late-fusion models on the clean test set. Tables 911 provide confidence, class-probability means, and branch-vs-fused MCC, respectively.
@@ Appendix F Full Branch Diagnostics for Late-Fusion Models
Model $\boldsymbol{\lambda}$ conf_a conf_b conf_f
Lite-CNN-Cons 0 $0.627352\pm 0.070257$ $0.554910\pm 0.020764$ $0.917221\pm 0.007478$
Lite-CNN-Cons 0.5 $0.508533\pm 0.004537$ $0.506999\pm 0.003842$ $0.913617\pm 0.005166$
Lite-CNN-Cons 1 $0.507229\pm 0.004763$ $0.507287\pm 0.005493$ $0.913873\pm 0.007414$
ResNet18P-Cons 0 $0.570899\pm 0.043481$ $0.567578\pm 0.040866$ $0.782159\pm 0.078004$
ResNet18P-Cons 0.5 $0.605562\pm 0.065867$ $0.611589\pm 0.066048$ $0.802972\pm 0.056561$
ResNet18P-Cons 1 $0.622989\pm 0.083976$ $0.624896\pm 0.082731$ $0.808066\pm 0.048449$
@@
Model $\boldsymbol{\lambda}$ pos_mean_a pos_mean_b pos_mean_f
Lite-CNN-Cons 0 $0.522772\pm 0.145785$ $0.511727\pm 0.058662$ $0.514203\pm 0.039776$
Lite-CNN-Cons 0.5 $0.502299\pm 0.009459$ $0.499700\pm 0.008371$ $0.511911\pm 0.056243$
Lite-CNN-Cons 1 $0.498687\pm 0.008587$ $0.498635\pm 0.009367$ $0.520245\pm 0.068696$
ResNet18P-Cons 0 $0.549286\pm 0.065936$ $0.513286\pm 0.074454$ $0.455765\pm 0.133288$
ResNet18P-Cons 0.5 $0.592447\pm 0.085277$ $0.590547\pm 0.095968$ $0.483531\pm 0.135293$
ResNet18P-Cons 1 $0.564726\pm 0.139656$ $0.564738\pm 0.140885$ $0.482701\pm 0.169748$

### v2

@@ Appendix A Implementation Details for Input Parameterization
Table˜3 summarizes the input parameterization and normalization settings.
Raw input
@@
Optional per-channel standardization
@@
Valid-range clipping
@@
Threat model defined in raw space
The threat model is always defined on $\mathbf{x}_{\mathrm{raw}}$ with an $\ell_{\infty}$ constraint $\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon_{\mathrm{adv}}$, where $\|\boldsymbol{\delta}\|_{\infty}=\max_{i}|\delta_{i}|$. When the model consumes standardized inputs, we enforce the same underlying raw-space budget by converting
@@
Item Setting Image storage 8-bit grayscale per view Load / scaling ToTensor() $\Rightarrow\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$ Normalization (use_norm) $\mathbf{x}_{\mathrm{norm}}=(\mathbf{x}_{\mathrm{raw}}-\mu)\oslash\sigma$ $\mu,\sigma$ estimation Randomly sample $\min(n_{\text{norm}},|{\rm train}|)$ training images without replacement (default $n_{\text{norm}}{=}512$); compute per-channel mean/std over $N{\times}H{\times}W$; clamp $\sigma\leftarrow\max(\sigma,10^{-6})$ Clipping (raw) $[0,1]$ Clipping (norm) $\big[(0-\mu)\oslash\sigma,(1-\mu)\oslash\sigma\big]$ Threat model space raw space; when attacking norm, use Eq. 13
Appendix B View Rendering and Multi-View Construction Details
This appendix details the view construction summarized in Section˜3. We follow the MODII-style design for multi-view financial image construction [11] and adopt standard time-series-to-image principles commonly used in financial prediction pipelines [4].
OHLCV rendering (ohlcv)
Each sample is rendered on a black canvas at a fixed resolution of $256\times 256$ pixels. The upper region (75% of the height) visualizes OHLC as candlestick-style glyphs (a high–low vertical stroke with open/close ticks). We shade the open–close range in gray to reflect the interval-valued view of price movements [29]. The lower region (25% of the height) encodes trading volume as a bar chart. Volumes are normalized by a rolling percentile scaler, i.e., divided by the trailing 95th percentile computed over a fixed reference window of length $L_{\mathrm{vol}}$, and then clipped to $[0,1]$.
Indicator-matrix rendering (indic)
We compute a fixed set of $N=15$ technical indicators over a lookback window of $L_{\mathrm{indic}}=15$ trading days. Concretely, we use RSI(14), WILLR(14), WMA(10), SMA(15), EMA(12), DEMA(14), TEMA(18), CCI(20), CMO(14), MACD(hist; 8,17,9), PPO(8,17; EMA), ROC(10), MFI(14), ADX(14), and SAR(0.02, 0.2); all choices and hyperparameters are fixed a priori. For each sample, we stack the most recent $L_{\mathrm{indic}}$ values into an $N\times L_{\mathrm{indic}}$ matrix (rows: indicators; columns: time) and apply per-row min–max scaling to $[0,255]$ within the current window. To keep the spatial layout consistent across splits, we compute a single indicator row order on the training split via average-linkage hierarchical clustering on the absolute Spearman correlation (distance $1-|\rho|$) with optimal leaf ordering (falling back to a spectral ordering if unavailable), and reuse it for validation/test. :contentReference[oaicite:0]index=0 :contentReference[oaicite:1]index=1 Finally, the $15\times 15$ matrix is upsampled with nearest-neighbor interpolation to $256\times 256$ (to match the image encoder input size) and rendered as a grayscale image.
Appendix C Model Architecture and Fusion Details
Table˜4 provides an implementation-level summary of all backbones and fusion heads used in the main experiments; the remainder of this appendix specifies layer-wise and fusion-head details for reproducibility.
[…]
Notation
@@ Appendix C Model Architecture and Fusion Details
Lite-CNN (PaperCNN)
Our Lite-CNN baseline is a compact CNN with three convolution blocks and two pooling operations, followed by a 2-layer MLP feature head and a final linear classifier. Concretely: (1) $\mathrm{Conv}_{3\times 3}(C\!\rightarrow\!16)$ + ReLU; (2) $\mathrm{Conv}_{3\times 3}(16\!\rightarrow\!16)$ + ReLU + $\mathrm{MaxPool}(2,2)$; (3) $\mathrm{Conv}_{3\times 3}(16\!\rightarrow\!32)$ + ReLU + $\mathrm{MaxPool}(3,3)$; then flatten and an MLP/classifier head $\mathrm{Linear}(\cdot\!\rightarrow\!256)\!\rightarrow\!\mathrm{ReLU}\!\rightarrow\!\mathrm{Dropout}\!\rightarrow\!\mathrm{Linear}(256\!\rightarrow\!128)\!\rightarrow\!\mathrm{ReLU}\!\rightarrow\!\mathrm{Linear}(128\!\rightarrow\!2)$. The flatten dimension is inferred automatically from the input resolution ($\times256$).
ResNet18-P
@@
Late fusion with optional consistency (*-Cons)
@@
Appendix D Additional Evaluation Details
From probabilities to labels
Reporting
@@ Appendix D Additional Evaluation Details
Filtering protocol for min_move
When using min_move ($\tau$), we restrict evaluation to the conditional subset defined by $|r_{t+1}|\geq\tau$. Because inclusion depends on the realized next-day return, min_move is an offline benchmark definition rather than an inference-time decision rule. To keep comparisons controlled across split variants and avoid inconsistencies near split boundaries, we apply the filter once on the full timeline prior to splitting, and then run the time-block split with embargo on the filtered index set (Section˜5.4, D). The resulting sample-size impact across $\tau$ values is summarized in Table˜5.
Time-Block Split and Embargo
This paragraph complements Section˜5.4 with implementation details of the time-block split with embargo [14, 28].
Let $B$ denote the block size (block_size) and $D_{\mathrm{emb}}$ the embargo length in trading days (embargo_days). We drop $K=\lceil D_{\mathrm{emb}}/B\rceil$ adjacent blocks around split boundaries to ensure a temporal gap of at least $D_{\mathrm{emb}}$ days (possibly larger due to block granularity) [14, 28]. Table˜5 reports, for each min_move setting, the aligned sample counts before splitting and the final counts retained after applying the embargoed time-block split, together with the resulting embargo percentage. All reported results use the same split protocol across models and settings.
min_move Aligned Train/Val/Test Used Embargo 0 4839 3420/630/699 4749 1.86% 0.002 3666 2580/480/516 3576 2.45% 0.004 2696 1890/330/386 2606 3.34% 0.006 1989 1380/240/279 1899 4.52% 0.008 1471 1050/150/181 1381 6.12% 0.010 1054 750/90/124 964 8.54%
Appendix E Training Protocol and Reproducibility Details
All trainable baselines are implemented in PyTorch and trained under a unified optimization and checkpoint-selection protocol unless stated otherwise.
Optimization and checkpoint selection
@@ Appendix E Training Protocol and Reproducibility Details
Data loading and leakage-avoiding normalization
All images are resized to $256\times 256$ and loaded in raw $[0,1]$ space. When enabled (use_norm), we apply per-channel standardization using $(\boldsymbol{\mu},\boldsymbol{\sigma})$ estimated on the training split only; see A. Two-view inputs follow the fixed view-to-channel mapping in Section˜3.3.
Early stopping
We train for at most 40 epochs with early stopping (patience=8). We use a model-dependent minimum training epoch to stabilize optimization: min_epoch=15 for Lite-CNN and Lite-CNN-cons, and min_epoch=2 for ResNet18-P and ResNet18-P-cons.
Adversarial evaluation
Adversarial evaluation is performed at test time only under the view-aligned protocol (Section˜4.2). Attack hyperparameters and budget conversion for standardized inputs are provided in F and A.
Reproducibility
Appendix F Attack Hyperparameters and Implementation Notes
Perturbation budgets
Unless stated otherwise, perturbation budgets are defined in raw $[0,1]$ pixel space under an $\ell_{\infty}$ constraint, following standard adversarial-example evaluations [18, 1, 2]. We evaluate $\epsilon_{\mathrm{adv}}\in E$ as defined in Eq. (5), where $\epsilon_{\mathrm{adv}}=0$ corresponds to clean evaluation. When sweeping budgets, we run the evaluation once per $\epsilon_{\mathrm{adv}}$. Table˜6 summarizes the attack settings.
Loss and attack objective
@@ Appendix F Attack Hyperparameters and Implementation Notes
$\max_{\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon_{\mathrm{adv}}}\ \mathcal{L}_{\mathrm{CE}}\!\left(f(\mathbf{x}+\boldsymbol{\delta}),y\right).$
@@
FGSM (one-step)
FGSM is implemented as a single-step update followed by clipping to the valid input range [1].
PGD (multi-step)
PGD performs $K=10$ iterative updates with step size $\alpha=\epsilon_{\mathrm{adv}}/K$; we use a deterministic initialization (random_start==False) from the clean input, and project after each step to enforce the $\ell_{\infty}$ constraint [2].
Table 6: Attack settings used in our experiments. Budgets are defined in raw $[0,1]$ space. When attacking standardized inputs, we convert $(\epsilon_{\mathrm{adv}},\alpha)$ using Eq. 13.
Attack Objective Budget Steps $K$ Step size $\alpha$ Random init. FGSM untargeted CE (maximize) $\epsilon_{\mathrm{adv}}\in E$ 1 $\alpha=\epsilon_{\mathrm{adv}}$ No PGD untargeted CE (maximize) $\epsilon_{\mathrm{adv}}\in E$ 10 $\alpha=\epsilon_{\mathrm{adv}}/10$ No
Appendix G Clean Baselines: Full Numerical Commentary
This appendix provides the full numerical commentary for Section˜6.1 (clean evaluation), without changing any reported values or conclusions. Representative values are summarized in Table˜1; full curves are shown in Figure˜2.
Minimum-movement filtering defines the evaluation regime
Figure˜2 shows that the minimum-movement threshold $\tau$ (Section˜5.3) largely determines whether the task is learnable on the ex-post filtered subsets defined by min_move. Without filtering or with weak filtering ($\tau\in\{0,0.002,0.004\}$), most deep models hover around zero MCC (sometimes negative), indicating near-chance behavior on the full set. As $\tau$ increases to $0.006$ and $0.008$, performance improves sharply, consistent with restricting evaluation to subsets that exclude near-zero moves and thus reduce direction-label ambiguity. Aggregating across all deep configurations (Lite-CNN/ResNet18-P and *-cons, all input modes), the mean MCC rises from 0.002-0.002 at $\tau=0$ to $0.082$ at $\tau=0.006$ and peaks at $0.092$ at $\tau=0.008$, before dropping to $0.030$ at $\tau=0.010$. The drop at $\tau=0.010$ aligns with the reduced effective sample size (Table˜5), which increases variance and destabilizes training. We stress that $\tau$ is used here to define evaluation regimes ex post, not as an inference-time filtering rule.
Lite-CNN: single-view is more robust under weak filtering; dual-view becomes beneficial once labels stabilize
Late fusion drives the main gain; consistency regularization yields modest, non-monotonic effects
ResNet18-P: early fusion underperforms; late fusion partially recovers but variance is higher
LogReg as a diagnostic baseline
Appendix H Full Branch Diagnostics for Late-Fusion Models
Branch diagnostics for late-fusion two-view models (clean test, $\tau=0.006$)
To interpret robustness trends for *-cons models, we inspect clean-test branch diagnostics that summarize cross-branch alignment (agree_ab, sym_kl with T=2T{=}2) and branch-wise versus fused MCC (Table˜7; full results in Tables 810). For Lite-CNN-Cons, enabling consistency ($\lambda>0$) sharply increases alignment (higher agree_ab, much lower sym_kl), yet branch-wise MCC remains near zero while fused-head MCC stays substantially higher. This indicates that predictive signal is expressed primarily through the feature-level late-fusion pathway rather than strong standalone branch classifiers. For ResNet18P-Cons, larger $\lambda$ likewise improves alignment and is associated with higher fused MCC, but variance remains notable and robustness failures under indicator-view attacks persist.
@@ Appendix H Full Branch Diagnostics for Late-Fusion Models
Model $\boldsymbol{\lambda}$ agree_ab sym_kl mcc_a / mcc_b mcc_fuse Lite-CNN-Cons 0 $0.275090\pm 0.244057$ $164898\pm 0.125751 0.0$$019377\pm 0.061519\ /\ 0.005906\pm 0.051856$ $0.127459\pm 0.035946$ Lite-CNN-Cons 0.5 $830197\pm 0.185333 0.0$$00195\pm 0.000098 0.0$$001488\pm 0.053262\ /\ -0.001753\pm 0.076580$ $0.129607\pm 0.026449$ Lite-CNN-Cons 1 $803315\pm 0.205572 0.0$$00106\pm 0.000029 0.0$$006285\pm 0.057660\ /\ -0.015335\pm 0.052686$ $0.120755\pm 0.029468$ ResNet18P-Cons 0 $439964\pm 0.247314 0.0$$74461\pm 0.065610 0.0$$14477\pm 0.047350\ /\ 0.000779\pm 0.038367 0.0$$048051\pm 0.056600$ ResNet18P-Cons 0.5 $920251\pm 0.196921 0.0$$06523\pm 0.003518 0.0$$06042\pm 0.012812\ /\ -0.002165\pm 0.015076 0.0$$067112\pm 0.031388$ ResNet18P-Cons 1 $931452\pm 0.163280 0.0$$03425\pm 0.001379 0.0$$02706\pm 0.028186\ /\ 0.009707\pm 0.019310 0.0$$095146\pm 0.057933$
Tables 810 provide confidence, class-probability means, and branch-vs-fused MCC, respectively.
@@
Model $\boldsymbol{\lambda}$ conf_a conf_b conf_f Lite-CNN-Cons 0 $0.627352\pm 0.070257$ $0.554910\pm 0.020764$ $0.917221\pm 0.007478$ Lite-CNN-Cons 0.5 $0.508533\pm 0.004537$ $0.506999\pm 0.003842$ $0.913617\pm 0.005166$ Lite-CNN-Cons 1 $0.507229\pm 0.004763$ $0.507287\pm 0.005493$ $0.913873\pm 0.007414$ ResNet18P-Cons 0 $0.570899\pm 0.043481$ $0.567578\pm 0.040866$ $0.782159\pm 0.078004$ ResNet18P-Cons 0.5 $0.605562\pm 0.065867$ $0.611589\pm 0.066048$ $0.802972\pm 0.056561$ ResNet18P-Cons 1 $0.622989\pm 0.083976$ $0.624896\pm 0.082731$ $0.808066\pm 0.048449$
@@
Model $\boldsymbol{\lambda}$ pos_mean_a pos_mean_b pos_mean_f Lite-CNN-Cons 0 $0.522772\pm 0.145785$ $0.511727\pm 0.058662$ $0.514203\pm 0.039776$ Lite-CNN-Cons 0.5 $0.502299\pm 0.009459$ $0.499700\pm 0.008371$ $0.511911\pm 0.056243$ Lite-CNN-Cons 1 $0.498687\pm 0.008587$ $0.498635\pm 0.009367$ $0.520245\pm 0.068696$ ResNet18P-Cons 0 $0.549286\pm 0.065936$ $0.513286\pm 0.074454$ $0.455765\pm 0.133288$ ResNet18P-Cons 0.5 $0.592447\pm 0.085277$ $0.590547\pm 0.095968$ $0.483531\pm 0.135293$ ResNet18P-Cons 1 $0.564726\pm 0.139656$ $0.564738\pm 0.140885$ $0.482701\pm 0.169748$
@@
Model $\boldsymbol{\lambda}$ mcc_a mcc_b mcc_fuse Lite-CNN-Cons 0 $019377\pm 0.061519 0.0$$005906\pm 0.051856$ $0.127459\pm 0.035946$ Lite-CNN-Cons 0.5 $001488\pm 0.053262 0.0$$001753\pm 0.076580$ $0.129607\pm 0.026449$ Lite-CNN-Cons 1 $006285\pm 0.057660 0.0$$015335\pm 0.052686$ $0.120755\pm 0.029468$ ResNet18P-Cons 0 $14477\pm 0.047350 0.000$$0779\pm 0.038367 0.0$$048051\pm 0.056600$ ResNet18P-Cons 0.5 $006042\pm 0.012812 0.0$$02165\pm 0.015076 0.0$$067112\pm 0.031388$ ResNet18P-Cons 1 $002706\pm 0.028186 0.0$$09707\pm 0.019310 0.0$$095146\pm 0.057933$
Appendix I Sensitivity to the Operating Point: Adversarial Robustness at $\tau=0.008$
This appendix reports adversarial robustness results under a larger min_move threshold $\tau=0.008$. Because changing $\tau$ alters the labeling rule (and typically class balance), absolute clean MCC values should not be compared across $\tau$; we therefore focus on within-$\tau$ degradation patterns.
Table 11: PGD robustness summary at $\tau=0.008$ (min_move). We report MCC at two representative budgets ($0.25/255$ and $1/255$), the worst-case drop $\Delta_{\mathrm{worst}}$ over nonzero budgets $\{0.25,0.5,0.75,1\}/255$ (Eq. (8)), and the stress-test MCC at $\epsilon_{\mathrm{adv}}=2/255$.
Model Input Attack (PGD) MCC@0 MCC@0.25/255 MCC@1/255 $\boldsymbol{\Delta_{\mathrm{worst}}}$ MCC@2/255 Lite-CNN ohlcv on ohlcv 0.091 0.042 -0.101 -0.192 -0.279 Lite-CNN indic on indic 0.100 0.035 -0.187 -0.287 -0.410 Lite-CNN both (early) on ohlcv (ch0) 0.107 0.047 -0.101 -0.209 -0.300 Lite-CNN both (early) on indic (ch1) 0.107 0.043 -0.101 -0.208 -0.272 Lite-CNN both (early) on both 0.107 -0.008 -0.285 -0.393 -0.561 Lite-CNN-Cons ($\lambda{=}0$) both (late) on ohlcv (ch0) 0.100 0.061 -0.048 -0.148 -0.198 Lite-CNN-Cons ($\lambda{=}0$) both (late) on indic (ch1) 0.100 0.087 0.058 -0.043 0.032 Lite-CNN-Cons ($\lambda{=}0$) both (late) on both 0.100 0.057 -0.078 -0.178 -0.249 Lite-CNN-Cons ($\lambda{=}0.5$) both (late) on ohlcv (ch0) 0.090 0.062 -0.051 -0.141 -0.179 Lite-CNN-Cons ($\lambda{=}0.5$) both (late) on indic (ch1) 0.090 0.084 0.063 -0.027 0.043 Lite-CNN-Cons ($\lambda{=}0.5$) both (late) on both 0.090 0.053 -0.080 -0.170 -0.246 Lite-CNN-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.103 0.059 -0.070 -0.173 -0.219 Lite-CNN-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.103 0.099 0.064 -0.039 0.018 Lite-CNN-Cons ($\lambda{=}1$) both (late) on both 0.103 0.042 -0.098 -0.201 -0.269 ResNet18-P ohlcv on ohlcv 0.109 -0.489 -0.977 -1.087 -0.999 ResNet18-P indic on indic 0.108 -1.000 -1.000 -1.108 -1.000 ResNet18-P both (early) on ohlcv (ch0) 0.043 -0.453 -0.962 -1.005 -1.000 ResNet18-P both (early) on indic (ch1) 0.043 -0.830 -1.000 -1.043 -1.000 ResNet18-P both (early) on both 0.043 -0.936 -1.000 -1.043 -1.000 ResNet18-P-Cons ($\lambda{=}0$) both (late) on ohlcv (ch0) 0.079 -0.240 -0.799 -0.878 -0.974 ResNet18-P-Cons ($\lambda{=}0$) both (late) on indic (ch1) 0.079 -0.927 -1.000 -1.079 -1.000 ResNet18-P-Cons ($\lambda{=}0$) both (late) on both 0.079 -0.966 -1.000 -1.079 -1.000 ResNet18-P-Cons ($\lambda{=}0.5$) both (late) on ohlcv (ch0) 0.083 -0.246 -0.791 -0.875 -0.986 ResNet18-P-Cons ($\lambda{=}0.5$) both (late) on indic (ch1) 0.083 -0.921 -1.000 -1.083 -1.000 ResNet18-P-Cons ($\lambda{=}0.5$) both (late) on both 0.083 -0.964 -1.000 -1.083 -1.000 ResNet18-P-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.094 -0.247 -0.795 -0.889 -0.974 ResNet18-P-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.094 -0.944 -1.000 -1.094 -1.000 ResNet18-P-Cons ($\lambda{=}1$) both (late) on both 0.094 -0.978 -1.000 -1.094 -1.000
[…]